# DMER-Ranker: Learning to Rank Emotion Descriptions in the Absence of Ground Truth

### v1

@@ Abstract
Descriptive Multimodal Emotion Recognition (DMER) is a newly proposed task that aims to describe a person’s emotional state using free-form natural language. Unlike traditional discriminative methods that rely on predefined emotion taxonomies, DMER provides greater flexibility in emotional expression, enabling fine-grained and interpretable emotion representations. However, this free-form prediction paradigm introduces significant challenges in evaluation. Existing methods either depend on ground-truth descriptions that require substantial manual effort or simplify the task by shifting the focus from evaluating descriptions to evaluating emotion labels. However, the former suffers from the labor-intensive collection of comprehensive descriptions, while the latter overlooks critical aspects such as emotional temporal dynamics, intensity, and uncertainty. To address these limitations, we propose DMER-Ranker, a novel evaluation strategy that reformulates the traditional “prediction-ground truth” comparison into the “prediction-prediction” comparison, eliminating the need for ground-truth descriptions. We then employ the Bradley-Terry algorithm to convert pairwise comparison results into model-level rankings. Additionally, we explore the possibility of automatic preference prediction and introduce DMER-Preference, the first preference dataset specifically designed for human emotions. Our work advances the field of DMER and lays the foundation for more intelligent human-computer interaction systems.
@@ 1 Introduction
Descriptive Multimodal Emotion Recognition (DMER) aims to use free-form natural language to describe a person’s emotional state [1]. Unlike traditional discriminative approaches that rely on predefined emotion taxonomies [2, 3], DMER adopts a generative paradigm, offering greater flexibility in emotion representation. This shift not only enables more fine-grained and interpretable emotion expression but also presents significant opportunities for advancing emotion-intelligent human-computer interaction technologies [1]. Figure 1 illustrates the task distinction between descriptive and discriminative MER.
@@
However, the use of free-form descriptions complicates evaluation. Current evaluation strategies can be broadly classified into two categories. The first approach leverages ground-truth emotion descriptions and utilizes large language models (LLMs), such as OpenAI GPT [4], to measure similarity between predicted and ground-truth descriptions [1]. Nevertheless, emotions are inherently linked to diverse multimodal cues, including facial expressions, (micro-)gestures, head movements, hand actions, and vocal tones [5, 6, 7]. Consequently, generating a comprehensive and accurate description of a person’s emotional state is inherently challenging. Moreover, obtaining high-quality ground-truth descriptions entails substantial annotation costs [1]. Given these challenges, researchers propose bypassing the evaluation of emotion descriptions and focusing solely on emotion labels [8, 9]. While this approach is more cost-effective, it neglects key aspects of emotion expression present in descriptions, such as temporal dynamics, intensity, and uncertainty. Furthermore, it fails to consider the existence and reasonableness of clue-based analysis, resulting in an incomplete assessment of emotion descriptions. Figure 2 summarizes the existing evaluation strategies for DMER.
@@
Considering the limitations of existing methods, we introduce a novel evaluation approach for DMER, called DMER-Ranker, which shifts the evaluation paradigm from the “prediction-ground truth” comparison to the “prediction-prediction” comparison. Specifically, we provide two descriptions for a video and ask human annotators to determine which description better captures the character’s emotional state. We then employ the Bradley-Terry algorithm [10] to transform these pairwise comparison results into model-level rankings. Compared to previous methods (see Figure 2), our approach eliminates the need for costly ground-truth descriptions. Furthermore, it directly assesses model performance based on descriptions, rather than relying solely on emotion labels. Therefore, this method provides a valuable complement to existing evaluation strategies.
Beyond the evaluation strategy, we also investigate whether the aforementioned human annotation process can be automated using Multimodal Large Language Models (MLLMs), thereby further reducing evaluation costs. Current research on LLMs-as-judges primarily focuses on pure-text or vision-language inputs [11, 12], with limited exploration in the domain of human emotions, which is characterized by multimodality, temporal dynamics, and subtle variations [13, 14]. To address this gap, we conduct a pioneering study to examine the feasibility of using MLLMs as judges for DMER. We first construct a preference dataset, DMER-Preference. To build this dataset, we provide two emotion descriptions for each video and ask three annotators to evaluate which description better aligns with the character’s emotional state. Only samples with unanimous agreement among the three annotators are included in the final dataset. To the best of our knowledge, this is the first preference dataset specifically designed for human emotions. We then perform a comprehensive evaluation of various MLLMs on the preference prediction task and explore strategies to improve their consistency with human annotations. The main contributions of this paper are summarized as follows:
  • (Ranking Strategy) This paper introduces DMER-Ranker, a novel ranking strategy specifically designed for DMER. In contrast to prior approaches, our method eliminates the reliance on ground-truth descriptions but can still focus on evaluating complete emotion descriptions rather than isolated emotion labels. This work represents a significant advancement and valuable complement to existing evaluation methodologies.
@@ 2 Ranking Strategy: DMER-Ranker
This paper proposes a ranking strategy that eliminates the need for ground-truth descriptions. The core idea is to transform the traditional “prediction-ground truth” comparison into the “prediction-prediction” comparison. Specifically, suppose there are $M$ models, and let $m_{i}$ denote one of them. Consider pairwise comparisons between model $m_{i}$ and model $m_{j}$ as an example. To ensure reliability, we conduct preference comparisons across multiple samples. For each sample $x^{k},k\in[1,K]$, we prompt model $m_{i}$ and model $m_{j}$ to generate emotion descriptions for $x^{k}$, resulting in $d^{k}_{i}$ and $d^{k}_{j}$, respectively. We then engage multiple annotators to assign preference scores to $d^{k}_{i}$ and $d^{k}_{j}$, and their majority voted results are denoted as $p^{k}_{ij}$. Figure 3 illustrates the annotation platform layout, where description 1 and description 2 correspond to $d^{k}_{i}$ and $d^{k}_{j}$, respectively.
@@
Finally, we use gradient descent to estimate the advantage scores for each model. Models with higher scores are ranked higher. Section 4.6 provides an example to illustrate this ranking process.
@@ 3 Automatic Preference Prediction
The ranking strategy in Section 2 involves human annotations. In this section, we investigate whether this annotation process can be automated using MLLMs, given their strong capabilities in multimodal content understanding and reasoning [15]. To answer this question, we first need to construct a high-quality preference dataset tailored for human emotions, named DMER-Preference. Leveraging this dataset, we then reveal the feasibility of employing MLLMs for automatic preference prediction.
@@ 3.1 Preference Dataset: DMER-Preference
To build the dataset, we first collect 1,368 video samples and provide two emotion descriptions for each. Then, we recruit three annotators to assess these descriptions in the context of the corresponding video content and determine which one more accurately reflects the character’s emotional state. To ensure annotation quality, all annotators must pass a preliminary test focused on emotion recognition. Finally, we retain only samples with unanimous agreement among the three annotators, resulting in a total of 574 samples. This strict selection strategy ensures the reliability of the preference labels.

### v2

@@ Abstract
With the recent success of Large Language Models (LLMs), Descriptive Multimodal Emotion Recognition (DMER) has garnered increasing attention, which aims to describe a person’s emotional state using free-form natural language. Unlike traditional discriminative methods that rely on predefined emotion taxonomies, DMER offers greater flexibility in emotional expression, enabling fine-grained and interpretable emotion representations. However, this free-form prediction paradigm exposes significant challenges in evaluation. Existing methods either depend on ground-truth descriptions that require substantial manual annotations or simplify the task by shifting the focus from evaluating descriptions to evaluating emotion labels. However, this simplification overlooks critical aspects such as emotional temporal dynamics, intensity, and uncertainty. To address these limitations, we draw inspiration from Reinforcement Learning from Human Feedback (RLHF) and propose DMER-Ranker, a novel evaluation strategy that reformulates the traditional “prediction-ground truth” comparison into the “prediction-prediction” comparison, eliminating the need for ground-truth descriptions. We then employ the Bradley-Terry algorithm to convert pairwise comparison results into model-level rankings. Additionally, we explore the possibility of automatic preference prediction and introduce DMER-Preference, the first preference dataset specifically designed for human emotions. Our work advances the field of DMER and lays the foundation for more intelligent human-computer interaction systems.
@@ 1 Introduction
Descriptive Multimodal Emotion Recognition (DMER) [1], also known as Explainable MER (EMER) [2], aims to use free-form natural language to describe emotional states. Unlike traditional discriminative methods that depend on predefined emotion taxonomies [3, 4], DMER adopts a generative paradigm, offering greater flexibility in emotion representation. This enables fine-grained and interpretable emotional expression, presenting significant opportunities for advancing emotion-intelligent human-computer interaction technologies [5]. Recent advancements in Multimodal Large Language Models (MLLMs), with their rich vocabularies and multimodal understanding capabilities, have made this task feasible. Figure 1 shows the distinctions between descriptive and discriminative MER.
@@
However, evaluating the quality of such open-ended descriptions remains a non-trivial task. Current evaluation strategies can be broadly classified into two categories. The first approach leverages ground-truth emotion descriptions and utilizes LLMs, such as OpenAI GPT [6], to measure the similarity between predicted and ground-truth descriptions [2]. Nevertheless, emotions are inherently linked to diverse multimodal cues, including facial expressions, (micro-)gestures, head movements, hand actions, and vocal tones [7, 8, 9]. Consequently, generating a comprehensive and accurate description of a person’s emotional state is inherently challenging. Moreover, obtaining high-quality ground-truth descriptions entails substantial annotation costs [2]. Given these challenges, researchers propose bypassing the evaluation of emotion descriptions and focusing solely on emotion labels [1, 10]. While this approach is more cost-effective, it neglects key aspects of emotional expression present in descriptions, such as temporal dynamics, intensity, and uncertainty. Furthermore, it fails to consider the existence and reasonableness of clue-based analysis, resulting in an incomplete assessment of emotion descriptions. Figure 2 summarizes the existing evaluation strategies for DMER.
@@
Drawing inspiration from Reinforcement Learning from Human Feedback (RLHF) [11], we transform the complex problem of designing indicators to measure semantic similarity with ground truth into a simpler problem of learning human preferences. Specifically, we propose DMER-Ranker, shifting the evaluation paradigm from the “prediction-ground truth” comparison to the “prediction-prediction” comparison. In our annotation process, we provide two descriptions for a video and ask human annotators to determine which description better captures the character’s emotional state. We then employ the Bradley-Terry algorithm [12] to transform these pairwise comparison results into model-level rankings. Compared to previous methods (see Figure 2), our approach addresses the need for costly ground-truth descriptions. Furthermore, it directly assesses model performance based on descriptions, rather than relying solely on emotion labels. Therefore, this method provides a valuable complement to existing evaluation strategies.
Beyond the evaluation strategy, we also investigate whether the aforementioned human annotation process can be automated using MLLMs, thereby further reducing evaluation costs. Current research on LLMs-as-judges primarily focuses on pure-text or vision-language inputs [13, 14], with limited exploration in the domain of human emotions, which is characterized by multimodality, temporal dynamics, and subtle variations [15, 16]. To address this gap, we conduct a pioneering study to examine the feasibility of using MLLMs as judges for DMER. We first construct a preference dataset, DMER-Preference. To build this dataset, we provide two emotion descriptions for each video and recruit multiple annotators to evaluate which description better aligns with the character’s emotional state. Only samples with unanimous agreement among the all annotators are included in the final dataset. To the best of our knowledge, this is the first preference dataset specifically designed for human emotions. We then perform a comprehensive evaluation of various MLLMs on the preference prediction task and explore strategies to improve their consistency with human annotations. The main contributions of this paper are summarized as follows:
  • (Ranking Strategy) This paper draws inspiration from RLHF and introduces DMER-Ranker, a novel ranking strategy specifically designed for DMER. In contrast to prior approaches, our method eliminates the reliance on ground-truth descriptions but can still focus on evaluating complete emotion descriptions rather than isolated emotion labels. This work represents a significant advancement and valuable complement to existing evaluation methodologies.
@@ 2 Ranking Strategy: DMER-Ranker
This paper draws inspiration from RLHF and proposes a ranking strategy that eliminates the need for ground-truth descriptions. The core idea is to transform the traditional “prediction-ground truth” comparison into the “prediction-prediction” comparison. Section 4.6 provides an example to illustrate this ranking process. Specifically, suppose there are $M$ models, and let $m_{i}$ denote one of them. Consider pairwise comparisons between model $m_{i}$ and model $m_{j}$ as an example. To ensure reliability, we conduct preference comparisons across multiple samples. For each sample $x^{k},k\in[1,K]$, we prompt model $m_{i}$ and model $m_{j}$ to generate emotion descriptions for $x^{k}$, resulting in $d^{k}_{i}$ and $d^{k}_{j}$, respectively. We then recruit multiple annotators to assign preference scores to $d^{k}_{i}$ and $d^{k}_{j}$, and their majority voted results are denoted as $p^{k}_{ij}$. Figure 3 illustrates the annotation platform layout, where description 1 and description 2 correspond to $d^{k}_{i}$ and $d^{k}_{j}$, respectively.
@@
Finally, we use gradient descent to estimate the advantage scores for each model. Models with higher scores are ranked higher.
@@ 3 Automatic Preference Prediction
The above ranking strategy (see Section 2) relies on human annotations. In this section, we investigate the possibility of automating human feedback entirely. Specifically, we first construct a high-quality preference dataset tailored for human emotions, named DMER-Preference. Leveraging this dataset, we then reveal the feasibility of employing MLLMs for automatic preference prediction.
@@ 3.1 Preference Dataset: DMER-Preference
To build the dataset, we first collect 1,368 video samples and provide two emotion descriptions for each. Then, we recruit multiple annotators to assess these descriptions in the context of the corresponding video content and determine which one more accurately reflects the character’s emotional state. To ensure annotation quality, all annotators must pass a preliminary test focused on emotion recognition. Finally, we retain only samples with unanimous agreement among the all annotators, resulting in a total of 574 samples. This strict selection strategy ensures the reliability of the preference labels.